### v2

# Enhancing the Performance of Aspect-Based Sentiment Analysis Systems

Abstract
Aspect-based sentiment analysis aims to predict sentiment polarity with fine granularity. While Graph Convolutional Networks (GCNs) are widely utilized for sentimental feature extraction, their naive application for syntactic feature extraction can compromise information preservation. This study introduces an innovative edge-enhanced GCN, named SentiSys, to navigate the syntactic graph while preserving intact feature information, leading to enhanced performance. Specifically,we first integrate a bidirectional long short-term memory (Bi-LSTM) network and a self-attention-based transformer. This combination facilitates effective text encoding, preventing the loss of information and predicting long dependency text. A bidirectional GCN (Bi-GCN) with message passing is then employed to encode relationships between entities. Additionally, unnecessary information is filtered out using an aspect-specific masking technique. To validate the effectiveness of our proposed model, we conduct extensive evaluation experiments and ablation studies on four benchmark datasets. The results consistently demonstrate improved performance in aspect-based sentiment analysis when employing SentiSys. This approach successfully addresses the challenges associated with syntactic feature extraction, highlighting its potential for advancing sentiment analysis methodologies.
keywords:
Performance Evaluation, Aspect-based Sentiment Analysis Systems, Graph Convolutional Networks
1 Introduction
Aspect-based sentiment analysis [1] aims to classify sentences in aspect granularity. The sentiments of each aspect can be different even in the same sentence. Aspect-level sentiment analysis has been widely used in real-world applications. Suppose there is such a review comment among the restaurant reviews: “Although the menu is limited, the friendly staff provided us with a nice night.” If we pay attention to the aspect of “menu”, the sentiment is negative (“limited”); when it comes to “staff”, we can extract the opposite sentiment, i.e., positive (“friendly”). Such multi-aspect can introduce multi-sentiments in the same sentence, resulting in difficulty of sentiment prediction. To solve such difficulty, aspect-based sentiment analysis is attracting attention to improve services [2].
Figure 1: An example of spurious associations of a customer review system.
Neural networks are widely used for aspect-based sentiment analysis because of their promising results, especially recurrent neural networks (RNNs) [3] and the Transformer [4] have demonstrated their sequence modeling capabilities. [5, 6, 7, 8, 9]. However, these models are not affordable to establish precise dependencies between word pairs and aspects. Suppose such a sentence: “The dish looks mediocre but tastes surprisingly wonderful!” Although the aspect “dish” is linked to two opposite sentiments, “looks mediocre” and “tastes surprisingly wonderful”, we can easily understand that the latter is more tightly linked to the “dish”. However, this judgment may be difficult for these models due to the closer distance of the former to the “dish”. Convolutional neural networks (CNNs) [10] are studied to make predictions based on phrases that express sentiments more accurately. However, phrases generally have longer dependencies on the subject. For example, “tastes surprisingly wonderful” may express a positive sentiment more accurately than just “wonderful”. CNNs are not affordable for such long dependencies because of the local receptive field limitation [11].
Graph convolutional networks (GCNs) [12, 13] can decompose sentences into a tree structure composed of entities. Such a tree structure constructs a more direct path between aspects and related words. Therefore, GCNs resolve the long-distance dependency problem and facilitate the learning of the representation of nodes and their positional relationships [14]. However, the general GCNs lack the edge information. GCNs cannot capture the syntax and semantic relationships due to the lack of edge information resulting in poor performance [15, 16]. Figure 1 shows examples that illustrate these complex relationships in customer reviews. The aspect “service” is associated with “never had”, and its sentiment is positive in past reviews. However, in a new customer review, its sentiment is not positive. The network cannot determine the sentiment by relying on the phrase “never had”.
This study is an extension of our previous work [17], introducing an innovative edge-enhanced graph convolutional network (GCN) named SentiSys, designed to elevate the effectiveness of aspect-based sentiment analysis systems. In our approach, we leverage bidirectional long short-term memory (Bi-LSTM) to extract node word embeddings, and Bi-GCN is employed to model word dependencies within a tree structure, thereby enhancing word representations. The integration of these components facilitates accurate association of sentiment words with the target aspect, resulting in a notable performance improvement. The robustness of the proposed SentiSys model is substantiated through extensive evaluation experiments, confirming its high-performance capabilities. The key contributions of this work are encapsulated in three primary aspects, providing a comprehensive enhancement to aspect-based sentiment analysis methodologies.
  • Architecture based on neural networks: The SentiSys incorporates a powerful combination of Bi-LSTM, a transformer, and an edge-enhanced Bi-GCN. This intricate setup leverages linguistic syntax to construct a graph, which, in turn, serves as the foundation for extracting and representing textual content through a sophisticated GCN. This approach underscores the importance of effectively harnessing these cutting-edge technologies to decode sentiment and meaning within the text, making it a notable advancement in sentiment analysis.
  • Effective semantic edges: Employing a dependency tree not only reduces the gap between aspects and their related words but also establishes a syntactic pathway within the text. Additionally, SentiSys takes into account the edge information within syntactic dependency trees. This valuable data enhances the syntactic connection between associated words and the specific target aspect.
  • Performance enhancement: The proposed methodology’s efficacy is confirmed through validation on four standard benchmark datasets. The empirical findings conclusively illustrate that the SentiSys model surpasses other baseline models across the majority of performance metrics.
2 Related Work
2.1 Aspect-based Sentiment Analysis Systems
Conventional sentiment analysis methodologies typically involve training classifiers utilizing features such as bag-of-words and sentiment dictionaries [18]. However, both rule-based [19] and statistical-based strategies [20] require handcrafted features, revealing the limitations of these naive sentiment analysis methods, particularly when confronted with the complex intricacies inherent in customer review analysis. These shortcomings include challenges in capturing nuanced sentiments related to specific aspects of interest.
Motivated by the inadequacies of traditional sentiment analysis, the research landscape has progressively shifted towards aspect-based sentiment analysis. Aspect-based sentiment analysis encompasses four key tasks [21]: aspect term sentiment analysis (ATSA) [22], aspect category sentiment analysis (ACSA) [23], aspect term extraction (ATE) [24], and aspect category extraction (ACE) [25]. ATSA involves training classifiers to predict the sentiment polarity associated with specified aspect terms, addressing the need for a more granular understanding of sentiment expressions within the context of specific elements mentioned in reviews [26]. ACSA, on the other hand, aims to predict sentiment polarity concerning predefined sentiment categories, providing insights into broader sentiment trends related to predetermined aspects [27]. ATE and ACE, conceptualized as entity extraction challenges, strive to solve the intricate problem of identifying and understanding aspects within reviews. ATE focuses on extracting precise aspect terms, such as the previously mentioned example of “dish,” aiming to capture specific entities of interest [28]. Conversely, ACE aims to identify and categorize broader aspects, like the “food” category, encompassing terms like “dish.”
In addressing these tasks, the motivation lies in achieving a more nuanced and targeted comprehension of sentiment expressions within diverse contexts. By refining the understanding of sentiments associated with specific aspects, aspect-based sentiment analysis contributes to enhanced accuracy and applicability in sentiment analysis practices [29]. However, it is essential to acknowledge potential limitations, such as the need for robust methods to handle ambiguous language and context, which opens avenues for further research in advancing the effectiveness of aspect-based sentiment analysis methodologies.
2.2 RNN-based Sentiment Analysis Systems
The recent remarkable advances in neural networks have attracted the attention of researchers for applications [30, 31, 32, 33, 34]. The models proposed in [35] for target-dependent sentiment classification leverage neural networks. Specifically, the LSTM architecture, originally designed for sequence modeling, proves effective for aspect-based sentiment analysis tasks in an autoregressive manner. Building upon the LSTM framework, this approach enhances the architecture by incorporating target-specific information. Referred to as target-dependent LSTM (TD-LSTM), this model skillfully captures the dynamic interplay between a target word and its contextual words, discerningly selecting pertinent segments of the context to deduce sentiment polarity towards the target. The model is amenable to end-to-end training through standard backpropagation, employing the cross-entropy error of supervised sentiment classification as its loss function. While the TD-LSTM approach enhances sentiment classification by considering target-specific information, the nuanced understanding of the interplay between a target word and its contextual words remains a complex task. This limitation may impact the model’s ability to capture subtle sentiment nuances and context-specific variations accurately.
Similar to TD-LSTM, ATAE-LSTM [5] is crafted to emphasize contextual relationships, utilizing a Bi-LSTM and an attention-based LSTM, respectively. While these models exhibit strengths, they encounter difficulties in capturing extensive contextual information, especially when crucial details are distant from the target. Recognizing this limitation, RAM [7] and TNet-LF [36] extend ATAE-LSTM [5] by introducing a multi-attention and Bi-attention mechanism, respectively. These enhancements significantly augment the models’ capability to consider broader contextual information, effectively addressing the challenge of handling long-range dependencies. However, TNet-LF cannot fully capture the intricacies of sentence structure and grammar, limiting its ability to discern nuanced syntactic relationships crucial for a comprehensive understanding of sentiment in text. Additionally, its reliance on semantic aspects may result in suboptimal performance when dealing with complex syntactic structures in natural language. Sentic LSTM [37] stands as a sophisticated system designed for targeted aspect-based sentiment analysis. The core innovation lies in the infusion of commonsense knowledge into an attentive LSTM network. RACL [38] is a sentiment analysis system based on relation-aware collaborative learning. This system tackles fundamental sentiment analysis tasks, encompassing aspect term extraction, opinion term extraction, and aspect-level sentiment classification. Nevertheless, the absence of detailed assessments and comparisons with existing models impedes a comprehensive understanding of RACL’s effectiveness in real-world scenarios.
Of significance in this domain is the Transformer [4], an innovative architecture explicitly designed to excel in capturing intricate long-range dependencies within textual data. The Transformer’s demonstrated effectiveness positions it as a compelling choice for tasks requiring a robust treatment of extensive contextual relationships.
2.3 GCN-based Sentiment Analysis Systems
Unlike RNN-based sentiment analysis systems, GCN-based systems seek to represent language data as graphs, where words or phrases are nodes and relationships between them are edges. This representation allows for a more nuanced understanding of contextual dependencies, capturing intricate connections that may be challenging for traditional RNNs [39].
Several notable advancements in aspect-based sentiment analysis using GCNs have been introduced. ASGCN [40] showcases remarkable proficiency in capturing sentiment polarity across various aspects in a text. This is achieved through its adept modeling of relationships between aspects and their respective words utilizing GCNs. However, the efficacy of ASGCN is contingent on precise dependency tree parsing. Instances of inaccurate parsing pose a potential challenge, introducing a risk of suboptimal modeling. This directly undermines the fidelity of relationships between aspects and words. Consequently, the overall performance of ASGCN is adversely affected, especially within the domain of aspect-based sentiment analysis. AEGCN [41] seamlessly integrates GCNs with attention mechanisms, specifically embracing multi-head interactive attention and multi-head self-attention. The prowess of AEGCN crucially hinges on the precise encoding of attention. Any deviations or inaccuracies in this intricate process pose a risk of suboptimal modeling, directly impacting the relationships within the graph. SK-GCN [42] combines a syntactic dependency tree and GCN to model the syntax relationship more precisely. However, SK-GCN encounters a significant limitation attributed to the graph structure of commonsense knowledge. Specifically, the model faces challenges arising from an insufficient graph structure when representing commonsense knowledge. This limitation can impact the model’s ability to effectively leverage commonsense knowledge for aspect-level sentiment analysis. AGCN [43] aggregates the feature information of neighboring nodes to make the node representations related to each other. However, the effectiveness of AGCN heavily depends on the accuracy and completeness of the graph representation. In scenarios where the graph structure does not accurately capture the relationships within the data, AGCN may struggle to generate meaningful and representative features, leading to suboptimal performance.
Unlike previous studies, we present a novel integration ofBi-LSTM, Transformer, and GCNs to advance aspect-based sentiment analysis. SentiSys strategically leverages the strengths of each component: Bi-LSTM serves as a proficient feature extractor, capturing intricate patterns and semantic nuances within the text. Meanwhile, the Transformer excels at capturing long-distance dependencies, allowing our model to grasp the contextual relationships crucial for comprehensive sentiment analysis. Introducing GCN with a dependency parsing layer further enriches the model’s capabilities by incorporating syntactic information, enabling a more fine-grained understanding of the syntax relations within the text. By combining these components, SentiSys aims to provide a holistic and nuanced perspective on aspect-based sentiment analysis. This innovative amalgamation addresses the limitations of prior studies, fostering a more comprehensive and accurate sentiment analysis methodology.
3 SentiSys Model Description
In this section, we mainly describe the proposed SentiSys in seven aspects. The word embedding layer creates low-dimensional representations for each token in a vocabulary; an Bi-LSTM aims to extract features of sentences; a Transformer model is used to extract contextual features from a global perspective; a dependency parsing layer creates a tree structure of context grammar; an Bi-GCN processes the generated tree structure; an aspect-specific masking layer is used to reduce redundancy and enhance the accuracy; the sentiment classification layer classifies the sentiment of a given sentence. Figure 2 shows an overview architecture of the proposed SentiSys for the aspect-based sentiment analysis.
Figure 2: The architectural overview of the proposed SentiSys for aspect-based sentiment analysis comprises the following key components: (a) The initial word embedding layer parses sentences and embeds words into a vector space, forming the foundation for further analysis. (b) The Bi-LSTM network enables the capture of contextual information in a sentence for a deeper understanding of word relationships. (c) The transformer layer with a self-attention mechanism can explore the interrelationships and global word features in lengthy text, enhancing contextual understanding. (d) The dependency parsing layer can recognize the dependencies among the words in a sentence and construct a dependency tree to help explain the grammatical structure. (e) The Bi-GCN propagates information among the nodes of the parsed dependency tree using message passing, thus effectively modeling word relationships. (f) The aspect-specific masking technique refines the hidden representation by separating the information related to the aspect of interest, which reduces redundancy and improves the accuracy of aspect-based sentiment analysis. (g) The final sentiment classification layer is dedicated to classifying the sentiment of a given sentence into predefined categories, typically categorizing it as positive, negative, or neutral.
3.1 Word Embedding Layer
To tokenize a sentence consisting of $n$ words with an $m$-word aspect, we represent it as a sequence denoted as ${\bf S}$ in the following manner:
${\bf S}=\left\{{\bf w}_{1},{\bf w}_{2},\cdot\cdot\cdot,{\bf w}_{\gamma},\cdot\cdot\cdot,{\bf w}_{\gamma+m},\cdot\cdot\cdot,{\bf w}_{n}\right\},$ (1)
where ${\bf w}_{t}\in{\bf S}$ and $\gamma$ indicate a word token and starting index of an aspect, respectively. Each word token is then transformed into a low-dimensional vector denoted as ${\bf e}_{t}\in\mathbb{R}^{d_{w}}$, where $d_{w}$ denotes the dimension of the word embedding.
3.2 Bi-LSTM Network
Word embedding embeds the words in the sentence into low-dimensional vectors, and these embedding vectors are used as inputs to the Bi-LSTM network [44]. The utilization of Bi-LSTM allows for the extraction of contextual information from sentences. As mentioned in [45], the Bi-LSTM network comprises two components: the forward LSTM and the backward LSTM. In the forward LSTM, the hidden vector $\overrightarrow{\bf h}^{LSTM}_{t}$ at the current time step is computed by combining the input embedding ${\bf e}_{t}$ with $\overrightarrow{\bf h}^{LSTM}_{t-1}$ from the previous time step. The backward LSTM is constructed by utilizing ${\bf e}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t+1}$ from the subsequent time step. Finally, the hidden vectors $\overrightarrow{\bf h}^{LSTM}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t}$ are concatenated to create the overall hidden vector ${\bf h}^{LSTM}_{t}$. This process is illustrated as follows:
$\overrightarrow{\bf h}^{LSTM}_{t}=\text{LSTM}({\bf e}_{t},\overrightarrow{\bf h}^{LSTM}_{t-1}),~{}\overrightarrow{\bf h}^{LSTM}_{t}\in\mathbb{R}^{d_{h}},$ (2)
$\overleftarrow{\bf h}^{LSTM}_{t}=\text{LSTM}({\bf e}_{t},\overleftarrow{\bf h}^{LSTM}_{t+1}),~{}\overrightarrow{\bf h}^{LSTM}_{t}\in\mathbb{R}^{d_{h}},$ (3)
${\bf h}^{LSTM}_{t}=\overrightarrow{\bf h}^{LSTM}_{t}\oplus\overleftarrow{\bf h}^{LSTM}_{t},~{}{\bf h}^{LSTM}_{t}\in\mathbb{R}^{2\times d_{h}},$ (4)
where $d_{h}$ represents the dimension of the hidden layer in the Bi-LSTM, and $\oplus$ is the operator used for concatenation.
3.3 Transformer Network
A transformer encoder, as described in [4], utilizes self-attention mechanisms to examine the relationships and global characteristics of words within lengthy sentences. To incorporate the sequential positions of a sentence, sinusoidal positional encoding functions are employed, as outlined below:
$\text{P}_{pos,2i}=\sin(\frac{pos}{10000^{\frac{2i}{d_{model}}}}),$ (5)
$\text{P}_{pos,2i+1}=\cos(\frac{pos}{10000^{\frac{2i}{d_{model}}}}),$ (6)
where $pos$ represents the position of a word within a sentence, while $i$ signifies the $i$th dimension of the embedding, with the embedding dimension denoted as $d_{model}$. Subsequently, the combined vector of the embedding and positional encoding for the input sentence serves as the input to the transformer. The transformer computes attention weights according to the following procedure:
$\text{Attention}({\bf Q,K,V})=\text{Softmax}(\frac{{\bf QK}^{T}}{\sqrt{d_{k}}}){\bf V},$ (7)
where ${\bf Q}$, ${\bf K}$ and ${\bf V}$ represent the query, key and value matrices with dimensions $d_{k}$, $d_{k}$ and $d_{v}$, respectively. Usually, a transformer employs a multi-head attention mechanism to project the query, key and value matrices $h$ times to provide improved internal contextual associations. The mathematical expression for this process is as follows:
[…]
$\text{Head}_{i}=\text{Attention}({\bf QW}_{Q},{\bf KW}_{K},{\bf VW}_{V}),$ (9)
where ${\bf W,W}_{Q},{\bf W}_{K}$, and ${\bf W}_{V}$ indicate weight matrices of the self-attention mechanism. Next, the transformer encoder applies two normalization layers to extract deep features of the context. Consider ${\bf Z}_{out}$ as the result matrix generated by the transformer encoder, and its formulation can be expressed as follows:
${\bf Z}_{out}=\text{Transformer}({\bf S}).$ (10)
Finally, the calculated global attention scores play a pivotal role in the ultimate sentiment classification layer.
3.4 Dependency-parsing Layer
The output hidden vector of the Bi-LSTM network, denoted as ${\bf h}_{t}$, can be used as an input to the dependency-parsing layer in the Bi-GCN, which helps to identify word dependencies in a sentence by constructing a dependency tree. Formally, the Bi-GCN initiates a matrix to record the edge information weights. Initially, it derives the adjacency matrix ${\bf A}\in\mathbb{R}^{n\times n}$ from the sentence’s dependency tree. The conventional form of this adjacency matrix can be expressed as follows:
${\bf A}_{ij}=\begin{cases}1,\quad i=j.\\ 1,\quad i\neq j,~{}i~{}\text{and}~{}j~{}\text{have dependencies}.\\ 0,\quad\text{otherwise}.\end{cases}$ (11)
GCN lacks the ability to capture rich edge information due to the discrete nature of the adjacency matrix ( comprising only 0s and 1s). We present an adjacency matrix that contains non-discrete edge information, which is formulated as follows:
${\bf A}_{ij}=\begin{cases}1,\quad i=j,\\ \text{SDI}(i,j),\quad i\neq j,~{}i~{}\text{and}~{}j~{}\text{have dependencies},\\ 0,\quad\text{otherwise}.\end{cases}$ (12)
Here, SDI is used to compute syntactic dependency information between the word $i$ and the word $j$. The computation is summarized below:
$\text{SDI}(i,j)=\frac{\text{Count}(\text{sd}(i,j))}{\text{Count}(\text{sd}(\cdot))},$ (13)
where $\text{sd}(i,j)$ and $\text{sd}(\cdot)$ represent the syntactic relationship between word pair $i$ and $j$ and the overall syntax of the dataset.
3.5 Bi-GCN Network
A GCN is a specialized variant of a Convolutional Neural Network (CNN) specifically tailored for encoding structured graph data. In the scenario of a textual graph consisting of $n$ words, we generate an adjacency matrix based on the syntactic matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ to capture information regarding syntactic dependencies. The representation of word $i$ at layer $l$ is symbolized as ${\bf h}^{l}_{i}$. This graphical structure is conceptually depicted as follows:
${\bf h}^{l}_{i}=\sigma(\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1}+\mathbf{b}^{l}).$ (14)
In the GCN, $\mathbf{W}^{l}$ serves as the weight for linear transformation, $\mathbf{b}^{l}$ acts as the bias, and $\sigma$ represents the nonlinear activation function. The GCN harnesses the sentence’s dependency tree to enforce syntactic constraints, facilitating the identification of descriptive words associated with a specific aspect, taking into account their syntactic proximity. This capability equips GCN to effectively address situations where words related to the sentiment of an aspect are not adjacent. As a result, GCN is a well-suited choice for conducting aspect-based sentiment analysis. The update of each word’s representation is executed through the utilization of Bi-GCN in as follows:
$\overrightarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1},$ (15)
$\overleftarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}{\tilde{\mathbf{A}}_{ij}}^{T}\mathbf{W}^{l}h_{j}^{l-1},$ (16)
where $\overrightarrow{\bf h}^{l}_{i}$ and $\overleftarrow{\bf h}^{l}_{i}$ indicate the outputs of the forward and backward hidden layers for the word $i$ in the $l$th layer. Then, the forward and backward representations can be concatenated by
$\tilde{\bf h}^{l}_{i}=\overrightarrow{\bf h}^{l}_{i}\oplus\overleftarrow{\bf h}^{l}_{i},$ (17)
and
${\bf h}^{l}_{i}=\text{ReLU}(\frac{\tilde{\bf h}^{l}_{i}}{d_{i}+1}\mathbf{W}^{l}+\mathbf{b}^{l}),$ (18)
where $d_{i}$ represents the degree of the $i$th token in the adjacency matrix. Note that the parameters $\mathbf{W}^{l}$ and $\mathbf{b}^{l}$ are trainable in the Bi-GCN network.
3.6 Aspect-Specific Masking
We utilize the masking technique to reduce redundancy in the Bi-GCN hidden representation by hiding aspect-independent text. To be more specific, this masking can be depicted as follows:
${\bf h}_{i}^{l}=0,~{}1\leq i<\gamma,~{}\gamma+m-1<t\leq n.$ (19)
The result from the zero mask layer consists of aspect-focused characteristics.
$\mathbf{H}_{mask}^{l}=\left\{{\bf 0},\cdots,{\bf h}_{\gamma}^{l},\cdots,{\bf h}_{\gamma+m-1}^{l},\cdots,{\bf 0}\right\}.$ (20)
By employing graph convolution, the feature vector $\mathbf{H}_{mask}^{l}$ adeptly captures contextual information surrounding specific aspects. This encompassing process considers both syntactic dependencies and extended multi-word relationships. Moreover, an attention mechanism is harnessed to produce a vector representation of the concealed layer. The calculation of attention weights is executed as follows:
$\beta_{i}=\sum_{j=1}^{n}{\bf h}_{j}^{T}{\bf h}_{j}^{l},$ (21)
$\alpha_{i}=\frac{\exp(\beta_{i})}{\sum_{j=1}^{n}\exp(\beta_{j}))}.$ (22)
Finally, the sentiment classification process combines the cumulative representation scores mentioned earlier with the attention scores generated by the transformer. This results in the following sentiment classification:
$\mathbf{res}_{out}=\sum_{i=1}^{n}\alpha_{i}{\bf h}_{i}+{\bf Z}_{out}.$ (23)
3.7 Sentiment Classification
The $\mathbf{r}_{out}$ computed in the previous step is used as input to the fully connected layer for sentiment classification. The formula is as follows:
$\mathbf{prob}=\text{softmax}(\mathbf{res}_{out}\mathbf{W}_{p}+\mathbf{b}).$ (24)
3.8 Training Phase
Utilizing both cross-entropy and $\mathit{L_{2}}$ regularization [46], the loss function for SentiSys is formulated as follows:
$\text{Loss}=-\sum_{\hat{\mathit{p}}\in C}\log\mathbf{prob_{\hat{\mathit{p}}}}+\lambda\left\|\Theta\right\|_{2},$ (25)
where $\mathit{p}$ signifies the label, and $C$ represents the dataset. $\mathbf{prob_{\hat{\mathit{p}}}}$ denotes the $\hat{\mathit{p}}$th element of $\mathbf{prob}$. $\Theta$ and $\lambda$ are the trainable parameters and the coefficient for $\mathit{L_{2}}$ regularization, respectively. To provide a clearer insight into the SentiSys model, Algorithm 1 illustrates the training and testing procedures of the model, facilitating a more comprehensive understanding.
// Read sample sentences from datasets.
1 Tokenize a sentence with Eq.(1).
2 Embed words in word embedding layer.
3 Analyze dependency information of the sentence.
// A Bi-LSTM layer.
4 Use a Bi-LSTM layer to capture the contexual features of the sample sentence with Eqs. (2),(3), and (4).
// A transformer layer.
5 Encode positions with Eqs. (5) and (6).
6 Employ a transformer encoder architecture to capture the global features within a sentence with Eqs. (7), (8), (9), and (10).
// Dependency-parsing layer.
7 Replace the weight matrix (11) with Eq. (12) and save more available edge information for the graph.
Compute SDI with Eq. (13). // A Bi-GCN submodule.
8 Use Bi-GCN to extract the features of the structure graph with Eqs. (14), (15), (16), (17), and (18).
// Aspect-specific masking.
9 Utilize aspect-specific masking techniques to reduce redundant information in the Bi-GCN with Eqs. (19), (20), (21), (22), and (23).
// Sentiment classification.
10 Apply fully connected layers to construct a classifier to categorize the sentiment with Eq. (24).
// Training loss.
11 Train SentiSys with Eq. (25).
// Testing phase.
Evaluate the performance of SentiSys on test datasets.
Algorithm 1 Algorithm for the proposed SentiSys.
4 Experiments
The efficacy of the SentiSys model is substantiated through experiments on four widely employed benchmark datasets. Furthermore, the influence of GCN layers on performance is deliberated. Lastly, ablation studies are conducted to examine syntactic dependency, the edge weight matrix (SDI), and bi-direction.
4.1 Datasets
Four commonly used benchmark datasets were employed to assess the performance of the SentiSys model. The Twitter dataset [47] encompasses 6,940 Twitter posts from users, with 6,248 data points for training and 692 for testing. Additionally, three datasets, namely Rest14 [48], Rest15 [49], and Rest16 [50] contain customer comments and restaurant rating information. Rest14 comprises 4,728 customer comments, with 3,608 for training and 1,120 for testing. Rest15 includes 1,746 reviews, with 1,204 for training and 542 for testing. Rest16, the 2016 benchmark dataset, consists of 2,364 reviews, of which 1,748 are used for training, and the remaining 616 are for testing. To enhance dataset quality, sentences with conflicting polarity or unclear meaning were removed from the Rest15 and Rest16 datasets. A comprehensive overview of the four datasets’ statistics is presented in Table 1.
Table 1: Comprehensive overview of the four datasets’ statistics.
Dataset Split Positive Neutral Negative
Twitter Train 1561 3127 1560
Test 173 346 173
Rest14 Train 2164 637 807
Test 728 196 196
Rest15 Train 912 36 256
Test 326 34 182
Rest16 Train 1240 69 439
Test 469 30 117
4.2 Experimental Setup
In the experiments, we initialize the word embedding layer using a pre-trained GloVe vectors [51] with a dimensionality of 300. The hidden layer of the Bi-LSTM also has a dimensionality of 300. The Bi-GCN of SentiSys contains 3 layers with parameters initialized using a uniform distribution. The optimizer is Adam [52], which has a learning rate of 0.001. We use $\mathit{L_{2}}$ regularization with a batch size of 32. The training process is up to 100 epochs.
Table 2: Results of the comparison of the performance of baseline models with that of SentiSys. The values in gray cells indicate the maximum values.
Model Twitter Rest14 Rest15 Rest16
Acc (%) $\uparrow$ F1 (%) $\uparrow$ Acc (%) $\uparrow$ F1 (%) $\uparrow$ Acc (%) $\uparrow$ F1 (%) $\uparrow$ Acc (%) $\uparrow$ F1 (%) $\uparrow$
SVM 63.40 63.30 80.16 - - - - -
LSTM 69.56 67.70 78.13 67.47 77.37 55.17 86.80 63.88
TD-LSTM 70.80 69.00 78.00 66.73 76.39 58.70 82.16 54.21
MemNet 71.48 69.90 79.61 69.64 77.31 58.28 85.44 65.99
AOA 72.30 70.20 79.97 70.42 78.17 57.02 87.50 66.21
IAN 72.50 70.81 79.26 70.09 78.54 52.65 84.74 55.21
TNet-LF 72.98 71.43 80.42 71.03 78.47 59.47 89.07 70.43
ASGCN 71.53 69.68 80.86 72.19 79.34 60.78 88.69 66.64
MGAN 72.54 70.81 81.25 71.94 79.36 57.26 87.06 62.29
Sentic LSTM 70.66 67.87 79.43 70.32 79.55 60.56 83.01 68.22
CMLA-ALSTM - - 77.46 63.87 81.03 54.79 - -
RACL - - 81.42 69.59 83.26 59.85 - -
AEGCN 73.16 71.82 81.04 71.32 79.95 60.87 87.39 68.22
SK-GCN 71.97 70.22 80.36 70.43 80.12 60.70 85.17 68.08
AGCN 73.98 72.48 80.02 71.02 80.07 62.70 87.98 65.78
SentiSys 72.44 70.67 81.70 73.63 79.34 61.99 88.80 68.96
4.3 Baselines for Comparison
We first evaluate the performance of SentiSys by comparing it to baseline models. These baseline models are described below:
  • SVM: The machine learning-based support vector machine (SVM) is widely used in SemEval 2014 Task 4. This model utilizes a conventional methodology for the extraction of essential features, making it a well-established approach in natural language processing and sentiment analysis tasks.
  • LSTM [35]: This methodology leverages the LSTM’s ability to capture and encode sequential information, with the focus on the hidden layer vector at the last time step, which is instrumental in making sentiment predictions.
  • TD-LSTM [35]: This specialized LSTM variant is designed to extract interaction features that capture the intricate relationship between the target sentiment, often a specific aspect or entity, and the entire contextual information.
  • MemNet [53]: MemNet improves the quality of sentence representations by efficiently integrating external textual information.
  • AOA [54]: The attention-over-attention (AOA) mechanism-based architecture plays a pivotal role in enhancing the classification process. AOA’s unique approach involves employing attention mechanisms not just once but in a layered fashion, enabling more refined and context-aware classification.
  • IAN [8]: The interactive attention networks (IAN) present a novel approach to information extraction. IAN employs a dual LSTM architecture to independently capture representations for aspect terms and contextual information.
  • TNet-LF [36]: The transformation network with local feedback (TNet-LF) introduces a forward context-preserving transformation mechanism. This innovative approach serves a dual purpose: to retain and amplify the valuable information within the context. TNet-LF achieves this by implementing a transformation process that not only safeguards but also enriches the contextual information, ultimately enhancing the overall understanding of the data.
  • ASGCN [40]: The pioneering aspect-specific graph convolutional network (ASGCN) marks a significant milestone in the field of sentiment analysis by harnessing the power of graph convolution, a novel approach to analyzing and understanding sentiment with a focus on specific aspects. By utilizing the capabilities of a GCN, ASGCN enhances the accuracy and effectiveness of sentiment analysis in the context of different aspects, making it a noteworthy advancement in this domain.
  • MGAN [9]: The Multi-grained attention network (MGAN) was introduced as an innovative approach with the primary objective of discerning and capturing the nuanced sentiments expressed in a given text, particularly in relation to specific aspects or entities.
  • Sentic LSTM [37]: Sentic LSTM incorporates a trio of key components to enhance its sentiment analysis capabilities. These components include an aspect-specific attention mechanism, a module for embedding commonsense knowledge, and an attentive LSTM.
  • CMLA-ALSTM [55]: This network incorporates multiple layers of attention mechanisms, with tensor operators implemented within each layer. The use of tensor operators at each layer enhances the network’s ability to capture and process complex relationships and patterns within the data. This design enables the network to perform intricate computations and feature extraction, making it a powerful tool for tasks that require advanced attention mechanisms and data analysis.
  • RACL [38]: Relation-aware collaborative learning (RACL) is a methodology that enables subtasks to synchronize their efforts within a network through the process of relational propagation. RACL’s fundamental concept is the establishment of connections among various subtasks, enabling them to effectively communicate and collaborate.
  • AEGCN [41]: Attention-based encoding GCN (AEGCN) leverages a multi-headed attention mechanism in conjunction with an enhanced GCN for sentiment analysis. This sophisticated approach employs multiple attention heads to focus on various aspects of the input data and combines them with an improved GCN, resulting in a more powerful method for sentiment analysis.
  • SK-GCN [42]: The syntax- and knowledge-based GCN (SK-GCN) harnesses the structure of syntactic dependency trees and the wealth of commonsense knowledge by integrating them with a GCN.
  • AGCN [43]: The aggregated GCN (AGCN) effectively utilizes node feature information through a process of iteratively enhancing the representation of individual nodes based on their local neighborhoods.
4.4 Evaluation Measures
We use two performance metrics, Accuracy (Acc) and Macro F1 (F1), to evaluate and validate the effectiveness of the proposed SentiSys. Acc and F1 are widely recognized in the field of multi-classification tasks to assess the predictive and classification capabilities of models. Concretely, Acc quantifies the ratio of correctly classified samples and thus directly measures the accuracy of the model. F1 takes into account both accuracy and recall, offering a comprehensive perspective on the model’s performance. Formally, these two evaluation measures are calculated as follows:
$\text{Acc}=\frac{TP+TN}{TP+TN+FP+FN},$
$\text{Precision}=\frac{TP}{TP+FP},$
$\text{Recall}=\frac{TP}{TP+FN},$
[…]
where TP, TN, FP, and FN refer to the correctly identified positive cases, correctly identified negative cases, incorrectly identified positive cases, and incorrectly identified negative cases.
4.5 Evaluation Results
Table 2 presents a comparative analysis of the Acc and F1 scores between the proposed SentiSys model and 15 other baseline models. Notably, the SentiSys model demonstrates superior performance across various metrics in the Twitter, Rest14, and Rest15 datasets. While the AGCN model achieved the highest Acc and F1 scores in the Twitter dataset, it did not exhibit the same level of effectiveness as the SentiSys model in the Rest14 and Rest16 datasets. Specifically, on the Rest14 dataset, the SentiSys model stands out, achieving significantly higher Acc and F1 scores compared to the other baseline models. In the Rest15 dataset, RACL achieved the highest Acc score, but its F1 score was relatively low at 59.85. When it comes to the Rest16 dataset, the Acc and F1 scores of TNet-LF were only marginally superior to those of the SentiSys model. This observation could be attributed to the critical role of information from both parent and child nodes, highlighting the significance of considering the entire graph structure rather than solely processing it as a directed graph. Additionally, the limited sentence grammar in the Twitter dataset may have constrained prediction results. In summary, the experimental results show that our proposed SentiSys enhances the performance of aspect-based sentiment analysis systems compared with other baseline models, especially for TNet-LF on the Rest16 dataset. Furthermore, SentiSys outperforms ASGCN on all four benchmark datasets, which highlights the impact of word dependence on enhancing the classification results.
4.6 Effect of GCN Layers
Figure 3: Change curve of Acc scores with the number of GCN layers.
Figure 4: Change curve of F1 scores with the number of GCN layers.
Figures 3 and 4 show the effects of different GCN layers on the Acc and F1 scores of the proposed SentiSys model on the Rest14 dataset. X and Y axes indicate the number of GCN layers and the corresponding Acc and F1 scores, respectively. The experimental results show that the Acc and F1 scores tended to increase and then decrease as the number of layers increased. The Acc and F1 scores achieved the maximum values with the number of GCN layers of 3. The scores are 0.818 and 0.731, respectively.
4.7 Ablation Studies
Table 3: Ablation studies on syntactical dependencies ($\boldsymbol{D}$), edge weights ($\boldsymbol{EW}$), and bidirectional ($\boldsymbol{Bi}$) structures for the SentiSys.
Dataset Measures $\boldsymbol{D}$ w/o $\boldsymbol{EW}$ w/o $\boldsymbol{Bi}$ w/o SentiSys
Twitter Acc (%) $\uparrow$ 71.53 71.39 71.80 72.40
F1 (%) $\uparrow$ 69.68 69.37 70.42 70.67
Rest14 Acc (%) $\uparrow$ 80.86 81.34 81.57 81.70
F1 (%) $\uparrow$ 72.19 71.70 73.35 73.63
Rest15 Acc (%) $\uparrow$ 79.34 79.15 80.44 79.34
F1 (%) $\uparrow$ 60.78 60.06 61.58 61.99
Rest16 Acc (%) $\uparrow$ 88.69 88.28 88.37 88.80
F1 (%) $\uparrow$ 66.64 67.03 69.20 68.96
    The values in gray cells indicate the maximum values.
The three critical parts of the proposed SentiSys model are the grammar dependency tree, the edge weight matrix, and the Bi-GCN. Ablation studies are conducted to verify the effects of each submodule on performance. Table 3 shows the Acc and F1 scores of the proposed SentiSys model without grammar dependency, with the discrete edge weight matrix, and without the bidirectional GCN on the four benchmark datasets. The results show that the grammar dependency technique, which leads to a significant improvement in the model performance. The SentiSys model with dependencies outperformed the results without dependencies on all four datasets. Similarly, the results with the SDI edge weight matrix are better than the traditional edge weight matrix with the discrete number representations (i.e., 0 and 1) on all four datasets. It indicates that the edge weight matrix with SDI improves the performance. Additionally, the SentiSys model with the addition of Bi-GCN outperformed the results without bi-direction on six of the eight measures. Overall, the syntactic dependency tree, the edge weight matrix, and Bi-GCN effectively improved the model performance.
5 Conclusion
This study enhanced the performance of aspect-based sentiment analysis systems, called SentiSys, by leveraging an edge-enhanced bidirectional GCN. The SentiSys model employed Bi-LSTM to acquire a deep understanding of the syntax and semantics within the text. The hidden layer features derived from Bi-LSTM are utilized for constructing syntactic dependency trees in the Bi-GCN. Furthermore, SentiSys incorporated a transformer encoder layer designed to extract global information from the text, enhancing SentiSys’s ability to capture nuanced features in lengthy textual data. The dependency parsing layer within Bi-GCN extracts the graph structure based on the text’s syntax, facilitating the learning of connections between aspects and context. To further refine the model, an aspect-specific masking layer is implemented, reducing redundant information in the GCN’s hidden layer and ultimately improving the accuracy of the aspect-based sentiment analysis system. Finally, SentiSys combined the output from Bi-GCN and transformer, both of which serve as inputs for the sentiment classification classifier.
The SentiSys model faces two primary limitations: Firstly, the computation of the SDI, a noteworthy contribution of this study, was performed statistically. In future work, we aim to implement trainable vectors within the network to enrich the features associated with each edge. Secondly, the use of the transformer for feature enhancement prompts a discussion about the most effective stage for merging global information into the model. In future work, we aim to overcome the first limitation by enhancing the SentiSys model through the integration of trainable vectors within the network. This improvement is anticipated to facilitate a more dynamic and context-specific computation of the SDI, thereby enriching the features associated with each edge. Addressing the second limitation, we will delve into optimal strategies for effectively integrating global information into the model. We plan to explore techniques that maximize the transformer’s impact at specific stages, ensuring its contributions are strategically incorporated to enhance overall classification accuracy.

### v3

@@
-Enhancing the Performance of Aspect-Based Sentiment Analysis Systems
+
 Abstract
-Aspect-based sentiment analysis aims to predict sentiment polarity with fine granularity.
+Aspect-based sentiment analysis predicts sentiment polarity with fine granularity.
 While graph convolutional networks (GCNs) are widely utilized for sentimental feature extraction, their naive application for syntactic feature extraction can compromise information preservation.
@@
 Additionally, unnecessary information is filtered out using an aspect-specific masking technique.
-To validate the effectiveness of our proposed model, we conduct extensive evaluation experiments and ablation studies on four benchmark datasets.
-The results consistently demonstrate improved performance in aspect-based sentiment analysis when employing SentiSys.
-This approach successfully addresses the challenges associated with syntactic feature extraction, highlighting its potential for advancing sentiment analysis methodologies.
+To validate the effectiveness of our proposed model, we conduct extensive evaluation experiments on four benchmark datasets.
+The experimental results demonstrate enhanced performance in aspect-based sentiment analysis with the use of SentiSys.
 Keywords:
-Performance Evaluation, Aspect-based Sentiment Analysis Systems, Graph Convolutional Networks
+Aspect-based sentiment analysis systems Graph convolutional networks
 1 Introduction
@@
 Such multi-aspect can introduce multi-sentiments in the same sentence, resulting in difficulty of sentiment prediction.
-To solve such difficulty, aspect-based sentiment analysis is attracting attention to improve services [2].
-Figure 1: An example of spurious associations of a customer review system.
-Neural networks are widely used for aspect-based sentiment analysis because of their promising results, especially recurrent neural networks (RNNs) [3] and the Transformer [4] have demonstrated their sequence modeling capabilities.
-[5, 6, 7, 8, 9].
+To solve such difficulty, aspect-based sentiment analysis is attracting attention [2].
+Neural networks are widely used for aspect-based sentiment analysis because of their promising results, especially recurrent neural networks (RNNs) [38] and the Transformer [18, 30] have demonstrated their sequence modeling capabilities.
 However, these models are not affordable to establish precise dependencies between word pairs and aspects.
@@
 However, this judgment may be difficult for these models due to the closer distance of the former to the “dish”.
-Convolutional neural networks (CNNs) [10] are studied to make predictions based on phrases that express sentiments more accurately.
+Convolutional neural networks (CNNs) [28, 36] are studied to make predictions based on phrases that express sentiments more accurately.
 However, phrases generally have longer dependencies on the subject.
@@
 However, the general GCNs lack the edge information.
-GCNs cannot capture the syntax and semantic relationships due to the lack of edge information resulting in poor performance [15, 16].
-Figure 1 shows examples that illustrate these complex relationships in customer reviews.
-The aspect “service” is associated with “never had”, and its sentiment is positive in past reviews.
-However, in a new customer review, its sentiment is not positive.
-The network cannot determine the sentiment by relying on the phrase “never had”.
-This study is an extension of our previous work [17], introducing an innovative edge-enhanced graph convolutional network (GCN) named SentiSys, designed to elevate the effectiveness of aspect-based sentiment analysis systems.
+GCNs cannot capture the syntax and semantic relationships due to the lack of edge information resulting in poor performance [4].
+This study introduces an innovative edge-enhanced graph convolutional network (GCN) named SentiSys, designed to elevate the effectiveness of aspect-based sentiment analysis systems.
 In our approach, we leverage bidirectional long short-term memory (Bi-LSTM) to extract node word embeddings, and Bi-GCN is employed to model word dependencies within a tree structure, thereby enhancing word representations.
 The integration of these components facilitates accurate association of sentiment words with the target aspect, resulting in a notable performance improvement.
-The robustness of the proposed SentiSys model is substantiated through extensive evaluation experiments, confirming its high-performance capabilities.
-The key contributions of this work are encapsulated in three primary aspects, providing a comprehensive enhancement to aspect-based sentiment analysis methodologies.
-  • Architecture based on neural networks: The SentiSys incorporates a powerful combination of Bi-LSTM, a transformer, and an edge-enhanced Bi-GCN.
-This intricate setup leverages linguistic syntax to construct a graph, which, in turn, serves as the foundation for extracting and representing textual content through a sophisticated GCN.
-This approach underscores the importance of effectively harnessing these cutting-edge technologies to decode sentiment and meaning within the text, making it a notable advancement in sentiment analysis.
-  • Effective semantic edges: Employing a dependency tree not only reduces the gap between aspects and their related words but also establishes a syntactic pathway within the text.
-Additionally, SentiSys takes into account the edge information within syntactic dependency trees.
-This valuable data enhances the syntactic connection between associated words and the specific target aspect.
+The robustness of the proposed SentiSys is substantiated through extensive evaluation, confirming its high-performance capabilities.
+The main contributions are as follows:
+  • Architecture based on neural networks: SentiSys integrates Bi-LSTM, a transformer, and an edge-enhanced Bi-GCN, leveraging linguistic syntax to construct a graph.
+This graph serves as the foundation for extracting and representing textual content through a sophisticated GCN.
+This approach highlights the significance of harnessing these advanced technologies to decode sentiment and meaning within text, marking a notable advancement in sentiment analysis.
+  • Effective semantic edges: Using a dependency tree minimizes the gap between aspects and related words while establishing a syntactic pathway within the text.
+SentiSys also considers edge information within syntactic dependency trees, enhancing syntactic connections between associated words and the target aspect.
   • Performance enhancement: The proposed methodology’s efficacy is confirmed through validation on four standard benchmark datasets.
@@
 2.1 Aspect-based Sentiment Analysis Systems
-Conventional sentiment analysis methodologies typically involve training classifiers utilizing features such as bag-of-words and sentiment dictionaries [18].
+Conventional sentiment analysis methodologies typically involve training classifiers utilizing features such as bag-of-words and sentiment dictionaries.
 However, both rule-based [8] and statistical-based strategies [13] require handcrafted features, revealing the limitations of these naive sentiment analysis methods, particularly when confronted with the complex intricacies inherent in customer review analysis.
-These shortcomings include challenges in capturing nuanced sentiments related to specific aspects of interest.
+These challenges arise in capturing nuanced sentiments regarding specific aspects.
 Motivated by the inadequacies of traditional sentiment analysis, the research landscape has progressively shifted towards aspect-based sentiment analysis.
-Aspect-based sentiment analysis encompasses four key tasks [21]: aspect term sentiment analysis (ATSA) [22], aspect category sentiment analysis (ACSA) [23], aspect term extraction (ATE) [24], and aspect category extraction (ACE) [25].
+Aspect-based sentiment analysis encompasses four key tasks [3]: aspect term sentiment analysis (ATSA), aspect category sentiment analysis (ACSA), aspect term extraction (ATE), and aspect category extraction (ACE).
 ATSA involves training classifiers to predict the sentiment polarity associated with specified aspect terms, addressing the need for a more granular understanding of sentiment expressions within the context of specific elements mentioned in reviews [24].
@@
 By refining the understanding of sentiments associated with specific aspects, aspect-based sentiment analysis contributes to enhanced accuracy and applicability in sentiment analysis practices [9].
-However, it is essential to acknowledge potential limitations, such as the need for robust methods to handle ambiguous language and context, which opens avenues for further research in advancing the effectiveness of aspect-based sentiment analysis methodologies.
+Recognizing limitations is crucial, particularly in developing robust methods to address ambiguous context, thereby prompting further research to enhance the efficacy of aspect-based sentiment analysis.
 2.2 RNN-based Sentiment Analysis Systems
-The recent remarkable advances in neural networks have attracted the attention of researchers for applications [30, 31, 32, 33, 34].
+The recent remarkable advances in neural networks have attracted the attention of researchers for applications [16, 17, 37].
 The models proposed in [29] for target-dependent sentiment classification leverage neural networks.
@@
 While the TD-LSTM approach enhances sentiment classification by considering target-specific information, the nuanced understanding of the interplay between a target word and its contextual words remains a complex task.
-This limitation may impact the model’s ability to capture subtle sentiment nuances and context-specific variations accurately.
+This limitation affects model accuracy in capturing subtle sentiment nuances and context-specific variations.
 Similar to TD-LSTM, ATAE-LSTM [32] is crafted to emphasize contextual relationships, utilizing a Bi-LSTM and an attention-based LSTM, respectively.
@@
 Nevertheless, the absence of detailed assessments and comparisons with existing models impedes a comprehensive understanding of RACL’s effectiveness in real-world scenarios.
-Of significance in this domain is the Transformer [4], an innovative architecture explicitly designed to excel in capturing intricate long-range dependencies within textual data.
-The Transformer’s demonstrated effectiveness positions it as a compelling choice for tasks requiring a robust treatment of extensive contextual relationships.
 2.3 GCN-based Sentiment Analysis Systems
 Unlike RNN-based sentiment analysis systems, GCN-based systems seek to represent language data as graphs, where words or phrases are nodes and relationships between them are edges.
-This representation allows for a more nuanced understanding of contextual dependencies, capturing intricate connections that may be challenging for traditional RNNs [39].
+This representation enables a nuanced understanding of contextual dependencies, capturing intricate connections challenging for traditional RNNs [31].
 Several notable advancements in aspect-based sentiment analysis using GCNs have been introduced.
@@
 Consequently, the overall performance of ASGCN is adversely affected, especially within the domain of aspect-based sentiment analysis.
-AEGCN [41] seamlessly integrates GCNs with attention mechanisms, specifically embracing multi-head interactive attention and multi-head self-attention.
-The prowess of AEGCN crucially hinges on the precise encoding of attention.
-Any deviations or inaccuracies in this intricate process pose a risk of suboptimal modeling, directly impacting the relationships within the graph.
-SK-GCN [42] combines a syntactic dependency tree and GCN to model the syntax relationship more precisely.
-However, SK-GCN encounters a significant limitation attributed to the graph structure of commonsense knowledge.
-Specifically, the model faces challenges arising from an insufficient graph structure when representing commonsense knowledge.
-This limitation can impact the model’s ability to effectively leverage commonsense knowledge for aspect-level sentiment analysis.
-AGCN [43] aggregates the feature information of neighboring nodes to make the node representations related to each other.
-However, the effectiveness of AGCN heavily depends on the accuracy and completeness of the graph representation.
-In scenarios where the graph structure does not accurately capture the relationships within the data, AGCN may struggle to generate meaningful and representative features, leading to suboptimal performance.
-Unlike previous studies, we present a novel integration ofBi-LSTM, Transformer, and GCNs to advance aspect-based sentiment analysis.
-SentiSys strategically leverages the strengths of each component: Bi-LSTM serves as a proficient feature extractor, capturing intricate patterns and semantic nuances within the text.
-Meanwhile, the Transformer excels at capturing long-distance dependencies, allowing our model to grasp the contextual relationships crucial for comprehensive sentiment analysis.
-Introducing GCN with a dependency parsing layer further enriches the model’s capabilities by incorporating syntactic information, enabling a more fine-grained understanding of the syntax relations within the text.
-By combining these components, SentiSys aims to provide a holistic and nuanced perspective on aspect-based sentiment analysis.
-This innovative amalgamation addresses the limitations of prior studies, fostering a more comprehensive and accurate sentiment analysis methodology.
-3 SentiSys Model Description
-In this section, we mainly describe the proposed SentiSys in seven aspects.
-The word embedding layer creates low-dimensional representations for each token in a vocabulary; an Bi-LSTM aims to extract features of sentences; a Transformer model is used to extract contextual features from a global perspective; a dependency parsing layer creates a tree structure of context grammar; an Bi-GCN processes the generated tree structure; an aspect-specific masking layer is used to reduce redundancy and enhance the accuracy; the sentiment classification layer classifies the sentiment of a given sentence.
-Figure 2 shows an overview architecture of the proposed SentiSys for the aspect-based sentiment analysis.
-Figure 2: The architectural overview of the proposed SentiSys for aspect-based sentiment analysis comprises the following key components: (a) The initial word embedding layer parses sentences and embeds words into a vector space, forming the foundation for further analysis.
-(b) The Bi-LSTM network enables the capture of contextual information in a sentence for a deeper understanding of word relationships.
-(c) The transformer layer with a self-attention mechanism can explore the interrelationships and global word features in lengthy text, enhancing contextual understanding.
-(d) The dependency parsing layer can recognize the dependencies among the words in a sentence and construct a dependency tree to help explain the grammatical structure.
-(e) The Bi-GCN propagates information among the nodes of the parsed dependency tree using message passing, thus effectively modeling word relationships.
-(f) The aspect-specific masking technique refines the hidden representation by separating the information related to the aspect of interest, which reduces redundancy and improves the accuracy of aspect-based sentiment analysis.
-(g) The final sentiment classification layer is dedicated to classifying the sentiment of a given sentence into predefined categories, typically categorizing it as positive, negative, or neutral.
+Unlike prior research, we introduce a novel fusion of Bi-LSTM, Transformer, and GCNs to advance aspect-based sentiment analysis.
+SentiSys strategically utilizes each component’s strengths: Bi-LSTM acts as an adept feature extractor, capturing intricate patterns and semantic nuances.
+Meanwhile, the Transformer excels at capturing long-range dependencies, enabling our model to understand crucial contextual relationships.
+The introduction of GCN with a dependency parsing layer enriches the model by incorporating syntactic information, facilitating a more nuanced understanding of text syntax.
+By combining these elements, SentiSys aims to offer a comprehensive perspective on aspect-based sentiment analysis, addressing previous limitations and fostering a more accurate sentiment analysis methodology.
+3 SentiSys
+Figure 1: SentiSys for aspect-based sentiment analysis includes the following components: (a) Initial word embedding layer for parsing and embedding words into vectors.
+(b) Bi-LSTM network captures contextual information for deeper word relationship understanding.
+(c) Transformer layer with self-attention mechanism explores global word features, enhancing contextual understanding.
+(d) Dependency parsing layer recognizes word dependencies, constructing a dependency tree to explain grammatical structure.
+(e) Bi-GCN propagates information among tree nodes, effectively modeling word relationships.
+(f) Aspect-specific masking refines hidden representation, reducing redundancy and improving accuracy.
+(g) Sentiment classification layer categorizes sentence sentiment into predefined categories, such as positive, negative, or neutral.
+SentiSys consists of a word embedding layer for token representation, a Bi-LSTM for sentence feature extraction, a Transformer for global contextual feature extraction, a dependency parsing layer for context grammar tree creation, a Bi-GCN for processing the tree structure, an aspect-specific masking layer to reduce redundancy and enhance accuracy, and a sentiment classification layer for sentence sentiment classification.
+Figure 1 illustrates the architecture of SentiSys for aspect-based sentiment analysis.
 3.1 Word Embedding Layer
-To tokenize a sentence consisting of $n$ words with an $m$-word aspect, we represent it as a sequence denoted as ${\bf S}$ in the following manner:
-${\bf S}=\left\{{\bf w}_{1},{\bf w}_{2},\cdot\cdot\cdot,{\bf w}_{\gamma},\cdot\cdot\cdot,{\bf w}_{\gamma+m},\cdot\cdot\cdot,{\bf w}_{n}\right\},$ (1)
-where ${\bf w}_{t}\in{\bf S}$ and $\gamma$ indicate a word token and starting index of an aspect, respectively.
+To tokenize a sentence consisting of $n$ words with an $m$-word aspect, we represent it as a sequence denoted as ${\bf S}=\left\{{\bf w}_{1},{\bf w}_{2},\cdot\cdot\cdot,{\bf w}_{\gamma},\cdot\cdot\cdot,{\bf w}_{\gamma+m},\cdot\cdot\cdot,{\bf w}_{n}\right\}$, where ${\bf w}_{t}\in{\bf S}$ and $\gamma$ indicate a word token and starting index of an aspect, respectively.
 Each word token is then transformed into a low-dimensional vector denoted as ${\bf e}_{t}\in\mathbb{R}^{d_{w}}$, where $d_{w}$ denotes the dimension of the word embedding.
@@
 Finally, the hidden vectors $\overrightarrow{\bf h}^{LSTM}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t}$ are concatenated to create the overall hidden vector ${\bf h}^{LSTM}_{t}$.
-This process is illustrated as follows:
 $\overrightarrow{\bf h}^{LSTM}_{t}=\text{LSTM}({\bf e}_{t},\overrightarrow{\bf h}^{LSTM}_{t-1}),~{}\overrightarrow{\bf h}^{LSTM}_{t}\in\mathbb{R}^{d_{h}},$ (1)
@@
 $\text{Attention}({\bf Q,K,V})=\text{Softmax}(\frac{{\bf QK}^{T}}{\sqrt{d_{k}}}){\bf V},$ (6)
-where ${\bf Q}$, ${\bf K}$ and ${\bf V}$ represent the query, key and value matrices with dimensions $d_{k}$, $d_{k}$ and $d_{v}$, respectively.
-Usually, a transformer employs a multi-head attention mechanism to project the query, key and value matrices $h$ times to provide improved internal contextual associations.
-The mathematical expression for this process is as follows:
 […]
 $\text{Head}_{i}=\text{Attention}({\bf QW}_{Q},{\bf KW}_{K},{\bf VW}_{V}),$ (8)
-where ${\bf W,W}_{Q},{\bf W}_{K}$, and ${\bf W}_{V}$ indicate weight matrices of the self-attention mechanism.
+where ${\bf Q}$, ${\bf K}$ and ${\bf V}$ represent the query, key and value matrices with dimensions $d_{k}$, $d_{k}$ and $d_{v}$.
+${\bf W,W}_{Q},{\bf W}_{K}$, and ${\bf W}_{V}$ indicate weight matrices of the self-attention mechanism.
 Next, the transformer encoder applies two normalization layers to extract deep features of the context.
@@
 Initially, it derives the adjacency matrix ${\bf A}\in\mathbb{R}^{n\times n}$ from the sentence’s dependency tree.
-The conventional form of this adjacency matrix can be expressed as follows:
-${\bf A}_{ij}=\begin{cases}1,\quad i=j.\\ 1,\quad i\neq j,~{}i~{}\text{and}~{}j~{}\text{have dependencies}.\\ 0,\quad\text{otherwise}.\end{cases}$ (11)
-GCN lacks the ability to capture rich edge information due to the discrete nature of the adjacency matrix ( comprising only 0s and 1s).
-We present an adjacency matrix that contains non-discrete edge information, which is formulated as follows:
+We present an adjacency matrix that contains non-discrete edge information, formulated as
 ${\bf A}_{ij}=\begin{cases}1,\quad i=j,\\ \text{SDI}(i,j),\quad i\neq j,~{}i~{}\text{and}~{}j~{}\text{have dependencies},\\ 0,\quad\text{otherwise}.\end{cases}$ (10)
-Here, SDI is used to compute syntactic dependency information between the word $i$ and the word $j$. The computation is summarized below:
-$\text{SDI}(i,j)=\frac{\text{Count}(\text{sd}(i,j))}{\text{Count}(\text{sd}(\cdot))},$ (13)
-where $\text{sd}(i,j)$ and $\text{sd}(\cdot)$ represent the syntactic relationship between word pair $i$ and $j$ and the overall syntax of the dataset.
+Here, SDI is used to compute syntactic dependency information between the word $i$ and the word $j$.
+The computation is calculated as $\text{SDI}(i,j)=\frac{\text{Count}(\text{sd}(i,j))}{\text{Count}(\text{sd}(\cdot))}$, where $\text{sd}(i,j)$ and $\text{sd}(\cdot)$ represent the syntactic relationship between word pair $i$ and $j$ and the overall syntax of the dataset.
 3.5 Bi-GCN Network
@@
 In the scenario of a textual graph consisting of $n$ words, we generate an adjacency matrix based on the syntactic matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ to capture information regarding syntactic dependencies.
-The representation of word $i$ at layer $l$ is symbolized as ${\bf h}^{l}_{i}$.
-This graphical structure is conceptually depicted as follows:
+Word $i$ at layer $l$ can be represented as ${\bf h}^{l}_{i}$.
+This graphical structure is depicted as
 ${\bf h}^{l}_{i}=\sigma(\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1}+\mathbf{b}^{l}).$ (11)
-In the GCN, $\mathbf{W}^{l}$ serves as the weight for linear transformation, $\mathbf{b}^{l}$ acts as the bias, and $\sigma$ represents the nonlinear activation function.
-The GCN harnesses the sentence’s dependency tree to enforce syntactic constraints, facilitating the identification of descriptive words associated with a specific aspect, taking into account their syntactic proximity.
-This capability equips GCN to effectively address situations where words related to the sentiment of an aspect are not adjacent.
-As a result, GCN is a well-suited choice for conducting aspect-based sentiment analysis.
-The update of each word’s representation is executed through the utilization of Bi-GCN in as follows:
-$\overrightarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1},$ (15)
-$\overleftarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}{\tilde{\mathbf{A}}_{ij}}^{T}\mathbf{W}^{l}h_{j}^{l-1},$ (16)
-where $\overrightarrow{\bf h}^{l}_{i}$ and $\overleftarrow{\bf h}^{l}_{i}$ indicate the outputs of the forward and backward hidden layers for the word $i$ in the $l$th layer.
-Then, the forward and backward representations can be concatenated by
-$\tilde{\bf h}^{l}_{i}=\overrightarrow{\bf h}^{l}_{i}\oplus\overleftarrow{\bf h}^{l}_{i},$ (17)
-and
-${\bf h}^{l}_{i}=\text{ReLU}(\frac{\tilde{\bf h}^{l}_{i}}{d_{i}+1}\mathbf{W}^{l}+\mathbf{b}^{l}),$ (18)
+In GCN, $\mathbf{W}^{l}$ is the weight for linear transformation, $\mathbf{b}^{l}$ acts as bias, and $\sigma$ is the nonlinear activation function.
+GCN utilizes the sentence’s dependency tree to enforce syntactic constraints, facilitating the identification of descriptive words associated with a specific aspect, considering their syntactic proximity.
+This capability enables GCN to effectively handle non-adjacent words related to aspect sentiment, making it a suitable choice for aspect-based sentiment analysis.
+The update of each word’s representation is executed through Bi-GCN as follows:
+$\overrightarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}\tilde{\mathbf{A}}_{ij}\mathbf{W}^{l}h_{j}^{l-1}~{}~{}\text{and}~{}~{}\overleftarrow{\bf h}^{l}_{i}=\sum_{j=1}^{n}{\tilde{\mathbf{A}}_{ij}}^{T}\mathbf{W}^{l}h_{j}^{l-1},$ (12)
+where $\overrightarrow{\bf h}^{l}_{i}$ and $\overleftarrow{\bf h}^{l}_{i}$ are the outputs of the forward and backward hidden layers for the word $i$ in the $l$th layer.
+The forward and backward representations are concatenated by
+$\tilde{\bf h}^{l}_{i}=\overrightarrow{\bf h}^{l}_{i}\oplus\overleftarrow{\bf h}^{l}_{i}~{}~{}\text{and}~{}~{}{\bf h}^{l}_{i}=\text{ReLU}(\frac{\tilde{\bf h}^{l}_{i}}{d_{i}+1}\mathbf{W}^{l}+\mathbf{b}^{l}),$ (13)
 where $d_{i}$ represents the degree of the $i$th token in the adjacency matrix. Note that the parameters $\mathbf{W}^{l}$ and $\mathbf{b}^{l}$ are trainable in the Bi-GCN network.
 3.6 Aspect-Specific Masking
-We utilize the masking technique to reduce redundancy in the Bi-GCN hidden representation by hiding aspect-independent text. To be more specific, this masking can be depicted as follows:
-${\bf h}_{i}^{l}=0,~{}1\leq i<\gamma,~{}\gamma+m-1<t\leq n.$ (19)
-The result from the zero mask layer consists of aspect-focused characteristics.
-$\mathbf{H}_{mask}^{l}=\left\{{\bf 0},\cdots,{\bf h}_{\gamma}^{l},\cdots,{\bf h}_{\gamma+m-1}^{l},\cdots,{\bf 0}\right\}.$ (20)
-By employing graph convolution, the feature vector $\mathbf{H}_{mask}^{l}$ adeptly captures contextual information surrounding specific aspects.
-This encompassing process considers both syntactic dependencies and extended multi-word relationships.
-Moreover, an attention mechanism is harnessed to produce a vector representation of the concealed layer.
-The calculation of attention weights is executed as follows:
-$\beta_{i}=\sum_{j=1}^{n}{\bf h}_{j}^{T}{\bf h}_{j}^{l},$ (21)
-$\alpha_{i}=\frac{\exp(\beta_{i})}{\sum_{j=1}^{n}\exp(\beta_{j}))}.$ (22)
-Finally, the sentiment classification process combines the cumulative representation scores mentioned earlier with the attention scores generated by the transformer.
-This results in the following sentiment classification:
+We utilize the masking technique to reduce redundancy in the Bi-GCN hidden representation by hiding aspect-independent text.
+This masking can be depicted as ${\bf h}_{i}^{l}=0,~{}1\leq i<\gamma,~{}\gamma+m-1<t\leq n$.
+The result from the zero mask layer consists of aspect-focused characteristics, represented as $\mathbf{H}_{mask}^{l}=\left\{{\bf 0},\cdots,{\bf h}_{\gamma}^{l},\cdots,{\bf h}_{\gamma+m-1}^{l},\cdots,{\bf 0}\right\}$.
+Utilizing graph convolution, $\mathbf{H}_{mask}^{l}$ adeptly captures contextual information, including syntactic dependencies and extended multi-word relationships.
+The attention mechanism for the hidden layer operates as follows:
+$\beta_{i}=\sum_{j=1}^{n}{\bf h}_{j}^{T}{\bf h}_{j}^{l}~{}~{}\text{and}~{}~{}\alpha_{i}=\frac{\exp(\beta_{i})}{\sum_{j=1}^{n}\exp(\beta_{j}))}.$ (14)
+Finally, the sentiment classification combines cumulative representation scores with attentions from the transformer, resulting in the sentiment classification:
 $\mathbf{res}_{out}=\sum_{i=1}^{n}\alpha_{i}{\bf h}_{i}+{\bf Z}_{out}.$ (15)
 3.7 Sentiment Classification
-The $\mathbf{r}_{out}$ computed in the previous step is used as input to the fully connected layer for sentiment classification. The formula is as follows:
-$\mathbf{prob}=\text{softmax}(\mathbf{res}_{out}\mathbf{W}_{p}+\mathbf{b}).$ (24)
-3.8 Training Phase
-Utilizing both cross-entropy and $\mathit{L_{2}}$ regularization [46], the loss function for SentiSys is formulated as follows:
+The computed $\mathbf{r}{out}$ from the previous step serves as input to the fully connected layer for sentiment classification, given by $\mathbf{prob}=\text{softmax}(\mathbf{r}{out}\mathbf{W}_{p}+\mathbf{b})$.
+Finally, utilizing cross-entropy and $\mathit{L_{2}}$ regularization [7], the loss function is as follows:
 $\text{Loss}=-\sum_{\hat{\mathit{p}}\in C}\log\mathbf{prob_{\hat{\mathit{p}}}}+\lambda\left\|\Theta\right\|_{2},$ (16)
@@
 $\mathbf{prob_{\hat{\mathit{p}}}}$ denotes the $\hat{\mathit{p}}$th element of $\mathbf{prob}$.
-$\Theta$ and $\lambda$ are the trainable parameters and the coefficient for $\mathit{L_{2}}$ regularization, respectively.
-To provide a clearer insight into the SentiSys model, Algorithm 1 illustrates the training and testing procedures of the model, facilitating a more comprehensive understanding.
-// Read sample sentences from datasets.
-1 Tokenize a sentence with Eq.(1).
-2 Embed words in word embedding layer.
-3 Analyze dependency information of the sentence.
-// A Bi-LSTM layer.
-4 Use a Bi-LSTM layer to capture the contexual features of the sample sentence with Eqs. (2),(3), and (4).
-// A transformer layer.
-5 Encode positions with Eqs. (5) and (6).
-6 Employ a transformer encoder architecture to capture the global features within a sentence with Eqs. (7), (8), (9), and (10).
-// Dependency-parsing layer.
-7 Replace the weight matrix (11) with Eq. (12) and save more available edge information for the graph.
-Compute SDI with Eq. (13). // A Bi-GCN submodule.
-8 Use Bi-GCN to extract the features of the structure graph with Eqs. (14), (15), (16), (17), and (18).
-// Aspect-specific masking.
-9 Utilize aspect-specific masking techniques to reduce redundant information in the Bi-GCN with Eqs. (19), (20), (21), (22), and (23).
-// Sentiment classification.
-10 Apply fully connected layers to construct a classifier to categorize the sentiment with Eq. (24).
-// Training loss.
-11 Train SentiSys with Eq. (25).
-// Testing phase.
-Evaluate the performance of SentiSys on test datasets.
-Algorithm 1 Algorithm for the proposed SentiSys.
+$\Theta$ and $\lambda$ are the trainable parameters.
 4 Experiments
-The efficacy of the SentiSys model is substantiated through experiments on four widely employed benchmark datasets.
-Furthermore, the influence of GCN layers on performance is deliberated.
-Lastly, ablation studies are conducted to examine syntactic dependency, the edge weight matrix (SDI), and bi-direction.
 4.1 Datasets
-Four commonly used benchmark datasets were employed to assess the performance of the SentiSys model.
-The Twitter dataset [47] encompasses 6,940 Twitter posts from users, with 6,248 data points for training and 692 for testing.
-Additionally, three datasets, namely Rest14 [48], Rest15 [49], and Rest16 [50] contain customer comments and restaurant rating information.
-Rest14 comprises 4,728 customer comments, with 3,608 for training and 1,120 for testing.
-Rest15 includes 1,746 reviews, with 1,204 for training and 542 for testing.
-Rest16, the 2016 benchmark dataset, consists of 2,364 reviews, of which 1,748 are used for training, and the remaining 616 are for testing.
-To enhance dataset quality, sentences with conflicting polarity or unclear meaning were removed from the Rest15 and Rest16 datasets.
-A comprehensive overview of the four datasets’ statistics is presented in Table 1.
-Table 1: Comprehensive overview of the four datasets’ statistics.
+Table 1: Statistics of the four datasets.
 Dataset Split Positive Neutral Negative
@@
 Test 469 30 117
+Four commonly used benchmark datasets were employed to assess the performance of the SentiSys model.
+The Twitter dataset [10] encompasses 6,940 Twitter posts from users, with 6,248 data points for training and 692 for testing.
+Additionally, three datasets, namely Rest14 [27], Rest15 [25], and Rest16 [26] contain customer comments and restaurant rating information.
+Rest14 comprises 4,728 customer comments, with 3,608 for training and 1,120 for testing.
+Rest15 includes 1,746 reviews, with 1,204 for training and 542 for testing.
+Rest16, the 2016 benchmark dataset, consists of 2,364 reviews, of which 1,748 are used for training, and the remaining 616 are for testing.
+To enhance dataset quality, sentences with conflicting polarity or unclear meaning were removed from the Rest15 and Rest16 datasets.
+Table 1 shows the statistics of the four datasets.
 4.2 Experimental Setup
-In the experiments, we initialize the word embedding layer using a pre-trained GloVe vectors [51] with a dimensionality of 300.
+We initialize the word embedding layer using a pre-trained GloVe vectors [23] with a dimensionality of 300.
 The hidden layer of the Bi-LSTM also has a dimensionality of 300.
@@
 The optimizer is Adam [39], which has a learning rate of 0.001.
-We use $\mathit{L_{2}}$ regularization with a batch size of 32.
-The training process is up to 100 epochs.
+We use $\mathit{L_{2}}$ regularization with a batch size of 32 for up to 100 epochs.
 Table 2: Results of the comparison of the performance of baseline models with that of SentiSys. The values in gray cells indicate the maximum values.
@@
 RACL - - 81.42 69.59 83.26 59.85 - -
-AEGCN 73.16 71.82 81.04 71.32 79.95 60.87 87.39 68.22
-SK-GCN 71.97 70.22 80.36 70.43 80.12 60.70 85.17 68.08
-AGCN 73.98 72.48 80.02 71.02 80.07 62.70 87.98 65.78
 SentiSys 72.44 70.67 81.70 73.63 79.34 61.99 88.80 68.96
-4.3 Baselines for Comparison
-We first evaluate the performance of SentiSys by comparing it to baseline models. These baseline models are described below:
-  • SVM: The machine learning-based support vector machine (SVM) is widely used in SemEval 2014 Task 4.
-This model utilizes a conventional methodology for the extraction of essential features, making it a well-established approach in natural language processing and sentiment analysis tasks.
-  • LSTM [35]: This methodology leverages the LSTM’s ability to capture and encode sequential information, with the focus on the hidden layer vector at the last time step, which is instrumental in making sentiment predictions.
-  • TD-LSTM [35]: This specialized LSTM variant is designed to extract interaction features that capture the intricate relationship between the target sentiment, often a specific aspect or entity, and the entire contextual information.
-  • MemNet [53]: MemNet improves the quality of sentence representations by efficiently integrating external textual information.
-  • AOA [54]: The attention-over-attention (AOA) mechanism-based architecture plays a pivotal role in enhancing the classification process.
-AOA’s unique approach involves employing attention mechanisms not just once but in a layered fashion, enabling more refined and context-aware classification.
-  • IAN [8]: The interactive attention networks (IAN) present a novel approach to information extraction.
-IAN employs a dual LSTM architecture to independently capture representations for aspect terms and contextual information.
-  • TNet-LF [36]: The transformation network with local feedback (TNet-LF) introduces a forward context-preserving transformation mechanism.
-This innovative approach serves a dual purpose: to retain and amplify the valuable information within the context.
-TNet-LF achieves this by implementing a transformation process that not only safeguards but also enriches the contextual information, ultimately enhancing the overall understanding of the data.
-  • ASGCN [40]: The pioneering aspect-specific graph convolutional network (ASGCN) marks a significant milestone in the field of sentiment analysis by harnessing the power of graph convolution, a novel approach to analyzing and understanding sentiment with a focus on specific aspects.
-By utilizing the capabilities of a GCN, ASGCN enhances the accuracy and effectiveness of sentiment analysis in the context of different aspects, making it a noteworthy advancement in this domain.
-  • MGAN [9]: The Multi-grained attention network (MGAN) was introduced as an innovative approach with the primary objective of discerning and capturing the nuanced sentiments expressed in a given text, particularly in relation to specific aspects or entities.
-  • Sentic LSTM [37]: Sentic LSTM incorporates a trio of key components to enhance its sentiment analysis capabilities.
-These components include an aspect-specific attention mechanism, a module for embedding commonsense knowledge, and an attentive LSTM.
-  • CMLA-ALSTM [55]: This network incorporates multiple layers of attention mechanisms, with tensor operators implemented within each layer.
-The use of tensor operators at each layer enhances the network’s ability to capture and process complex relationships and patterns within the data.
-This design enables the network to perform intricate computations and feature extraction, making it a powerful tool for tasks that require advanced attention mechanisms and data analysis.
-  • RACL [38]: Relation-aware collaborative learning (RACL) is a methodology that enables subtasks to synchronize their efforts within a network through the process of relational propagation.
-RACL’s fundamental concept is the establishment of connections among various subtasks, enabling them to effectively communicate and collaborate.
-  • AEGCN [41]: Attention-based encoding GCN (AEGCN) leverages a multi-headed attention mechanism in conjunction with an enhanced GCN for sentiment analysis.
-This sophisticated approach employs multiple attention heads to focus on various aspects of the input data and combines them with an improved GCN, resulting in a more powerful method for sentiment analysis.
-  • SK-GCN [42]: The syntax- and knowledge-based GCN (SK-GCN) harnesses the structure of syntactic dependency trees and the wealth of commonsense knowledge by integrating them with a GCN.
-  • AGCN [43]: The aggregated GCN (AGCN) effectively utilizes node feature information through a process of iteratively enhancing the representation of individual nodes based on their local neighborhoods.
 4.3 Evaluation Measures
 We use two performance metrics, Accuracy (Acc) and Macro F1 (F1), to evaluate and validate the effectiveness of the proposed SentiSys.
-Acc and F1 are widely recognized in the field of multi-classification tasks to assess the predictive and classification capabilities of models.
+Acc and F1 are widely recognized in the field of multi-classification tasks to assess the predictive and classification capabilities of models [35].
 Concretely, Acc quantifies the ratio of correctly classified samples and thus directly measures the accuracy of the model.
 F1 takes into account both accuracy and recall, offering a comprehensive perspective on the model’s performance.
-Formally, these two evaluation measures are calculated as follows:
-$\text{Acc}=\frac{TP+TN}{TP+TN+FP+FN},$
-$\text{Precision}=\frac{TP}{TP+FP},$
-$\text{Recall}=\frac{TP}{TP+FN},$
-[…]
-where TP, TN, FP, and FN refer to the correctly identified positive cases, correctly identified negative cases, incorrectly identified positive cases, and incorrectly identified negative cases.
 4.4 Evaluation Results
@@
 Notably, the SentiSys model demonstrates superior performance across various metrics in the Twitter, Rest14, and Rest15 datasets.
-While the AGCN model achieved the highest Acc and F1 scores in the Twitter dataset, it did not exhibit the same level of effectiveness as the SentiSys model in the Rest14 and Rest16 datasets.
+While the TNet-LF model achieved the highest Acc and F1 scores in the Twitter and Rest16 datasets, it did not exhibit the same level of effectiveness as the SentiSys model in the Rest14 and Rest15 datasets.
 Specifically, on the Rest14 dataset, the SentiSys model stands out, achieving significantly higher Acc and F1 scores compared to the other baseline models.
@@
 Figure 3: Change curve of F1 scores with the number of GCN layers.
-Figures 3 and 4 show the effects of different GCN layers on the Acc and F1 scores of the proposed SentiSys model on the Rest14 dataset.
+Figures 3 and 3 show the effects of different GCN layers on the Acc and F1 scores of the proposed SentiSys model on the Rest14 dataset.
 X and Y axes indicate the number of GCN layers and the corresponding Acc and F1 scores, respectively.
@@
 The scores are 0.818 and 0.731, respectively.
-4.7 Ablation Studies
-Table 3: Ablation studies on syntactical dependencies ($\boldsymbol{D}$), edge weights ($\boldsymbol{EW}$), and bidirectional ($\boldsymbol{Bi}$) structures for the SentiSys.
-Dataset Measures $\boldsymbol{D}$ w/o $\boldsymbol{EW}$ w/o $\boldsymbol{Bi}$ w/o SentiSys
-Twitter Acc (%) $\uparrow$ 71.53 71.39 71.80 72.40
-F1 (%) $\uparrow$ 69.68 69.37 70.42 70.67
-Rest14 Acc (%) $\uparrow$ 80.86 81.34 81.57 81.70
-F1 (%) $\uparrow$ 72.19 71.70 73.35 73.63
-Rest15 Acc (%) $\uparrow$ 79.34 79.15 80.44 79.34
-F1 (%) $\uparrow$ 60.78 60.06 61.58 61.99
-Rest16 Acc (%) $\uparrow$ 88.69 88.28 88.37 88.80
-F1 (%) $\uparrow$ 66.64 67.03 69.20 68.96
-    The values in gray cells indicate the maximum values.
-The three critical parts of the proposed SentiSys model are the grammar dependency tree, the edge weight matrix, and the Bi-GCN.
-Ablation studies are conducted to verify the effects of each submodule on performance.
-Table 3 shows the Acc and F1 scores of the proposed SentiSys model without grammar dependency, with the discrete edge weight matrix, and without the bidirectional GCN on the four benchmark datasets.
-The results show that the grammar dependency technique, which leads to a significant improvement in the model performance.
-The SentiSys model with dependencies outperformed the results without dependencies on all four datasets.
-Similarly, the results with the SDI edge weight matrix are better than the traditional edge weight matrix with the discrete number representations (i.e., 0 and 1) on all four datasets.
-It indicates that the edge weight matrix with SDI improves the performance.
-Additionally, the SentiSys model with the addition of Bi-GCN outperformed the results without bi-direction on six of the eight measures.
-Overall, the syntactic dependency tree, the edge weight matrix, and Bi-GCN effectively improved the model performance.
 5 Conclusion
-This study enhanced the performance of aspect-based sentiment analysis systems, called SentiSys, by leveraging an edge-enhanced bidirectional GCN.
-The SentiSys model employed Bi-LSTM to acquire a deep understanding of the syntax and semantics within the text.
-The hidden layer features derived from Bi-LSTM are utilized for constructing syntactic dependency trees in the Bi-GCN.
-Furthermore, SentiSys incorporated a transformer encoder layer designed to extract global information from the text, enhancing SentiSys’s ability to capture nuanced features in lengthy textual data.
-The dependency parsing layer within Bi-GCN extracts the graph structure based on the text’s syntax, facilitating the learning of connections between aspects and context.
-To further refine the model, an aspect-specific masking layer is implemented, reducing redundant information in the GCN’s hidden layer and ultimately improving the accuracy of the aspect-based sentiment analysis system.
-Finally, SentiSys combined the output from Bi-GCN and transformer, both of which serve as inputs for the sentiment classification classifier.
-The SentiSys model faces two primary limitations: Firstly, the computation of the SDI, a noteworthy contribution of this study, was performed statistically.
-In future work, we aim to implement trainable vectors within the network to enrich the features associated with each edge.
-Secondly, the use of the transformer for feature enhancement prompts a discussion about the most effective stage for merging global information into the model.
-In future work, we aim to overcome the first limitation by enhancing the SentiSys model through the integration of trainable vectors within the network.
-This improvement is anticipated to facilitate a more dynamic and context-specific computation of the SDI, thereby enriching the features associated with each edge.
-Addressing the second limitation, we will delve into optimal strategies for effectively integrating global information into the model.
-We plan to explore techniques that maximize the transformer’s impact at specific stages, ensuring its contributions are strategically incorporated to enhance overall classification accuracy.
+This study improved aspect-based sentiment analysis with SentiSys, using an edge-enhanced bidirectional GCN.
+Bi-LSTM provided deep text understanding, with hidden features used to construct syntactic dependency trees in Bi-GCN.
+SentiSys included a transformer encoder layer for global information extraction, enhancing feature capturing in lengthy text.
+The dependency parsing layer within Bi-GCN extracted graph structure based on text syntax, aiding aspect-context connections learning.
+An aspect-specific masking layer refined the model, reducing redundant information in GCN’s hidden layer and enhancing accuracy.
+Finally, SentiSys combined Bi-GCN and transformer outputs for sentiment classification.
+Two primary limitations were identified: statistical computation of SDI and transformer’s feature enhancement stage.
+Future work aims to address these limitations by integrating trainable vectors for dynamic SDI computation and exploring optimal strategies for transformer integration to enhance classification accuracy.
+Acknowledgements
+This research was supported by the National Natural Science Foundation of China (Grant No.
+62402022), the International Research Fellow of Japan Society for the Promotion of Science (Postdoctoral Fellowships for Research in Japan [Standard]), and KAKENHI (20K11830) Japan.